• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Deur

    Log entry time 16:01:55 on May 8,2010

    Entry number 321754

    keyword=Day shift summary

    Starting shift with production run 3632 on-going on Lead-Diamond target #2
    08:27: End of run 3632. When we tried to start run 3633, start run failed. We kill coda and retried. start run 3634 failed: roc31 not reporting. Rebooted roc31. Start coda again. Run started fine but we don't get any events. We reboot all rocs except roc23. It seems to be the cause (cannot get the prompt on roc23 to reboot it). Call rc. He will try to reboot it from home. No luck. He suggests to call a dithering expert. Physically rebooted the Happex crate. This seems to have fixed the problem.

    09:09 Start run 3636.
    09:23 restart dithering because we rebooted roc23. Dithering parameters are what they should be.
    Although I started the dithering, it seems that it was not doing anything during run 3636. So there may be no dithering during run 3636.
    10:00 Stop run 3636. Start run 3637 (we checked that dithering is now working, see Franco Garibaldi's entry).
    10:52 Stop run 3637. Start run 3638. ROC31 not responding. Since it's the same symptons as previously, I rebooted all the rocs (again can't get roc23 prompt so I did a hard reboot). Restarted coda. Restarted the dithering.
    11:08 Start run 3640. Seems the dithering doesn't work. Restarted it. Seems it is still not working. The RC is looking into it.
    11:21 MCC called. They have problems. Beam will be off for 10min or more.
    11:36: beam is back
    11:59 End of run 3640. RC ask us to wait before to start new run in order to fix left HSR Q1.
    12:09 Start new run 3641
    13:01 Run 3641 over. Switching helicty board frequency to 120Hz and T_stable=8233s
    13:07 Start pedestal calibration run 3642 CODA crashed 3642->44 are junk
    13:27 Ask MCC to turn of autogain for BPM. Start test run 3645
    13:30 Start BPM callibration run 3646. Dithering and feedback are off.
    Call MCC to ask for ramp down the beam (twice).
    14:01 Stop run 3646. Call MCC for beam off. Move target to empty position. Ramping down septum magnet. Ask for controll access for Compton work.
    14:24 MCC called to tell us that the BPM autogain are being turned back on.
    14:54 Per MCC request, we turned off dithering.